OpenStudy (anonymous):

What fraction is equivalent to 7 and 5/9?

ganeshie8 (ganeshie8):

What fraction is... there are infinitely many fractions that equal the given fraction

ganeshie8 (ganeshie8):

you need to look at options and pick the one that matches

OpenStudy (anonymous):

There has to be a following?

ganeshie8 (ganeshie8):

So provide options for questions like these okay ?

OpenStudy (anonymous):

A. 41/9 B. 68/9 C. 75/9 D. 75/3

OpenStudy (anonymous):

here they are :)

OpenStudy (anonymous):

 Do you know how to add fractions?

ganeshie8 (ganeshie8):

7 `and` 5/9 replace and with `+`

ganeshie8 (ganeshie8):

get a common denominator and add the numerators : \(\large 7 + \dfrac{5}{9}\) \(\large \dfrac{7\times 9}{9} + \dfrac{5}{9}\) \(\large \dfrac{7\times 9 + 5}{9} \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

 I just want to point that out, you can imagine if there's a 1 in the denominator so it looks like a "actual fraction."